The equation of a transverse wave traveling along a very long string is y = 3.96 sin(0.0444πx+ 7.89πt), where x and y are expressed in centimeters and t is in seconds. Determine (a) the amplitude, (b) the wavelength, (c) the frequency, (d) the speed, (e) the direction of propagation of the wave and (f) the maximum transverse speed of a particle in the string. (g) What is the transverse displacement at x = 1.05 cm when t = 0.843 s?
Answer #1

Given equation: y(x,t)= 3.96sin(0.0444\pi x+7.89\pi t)

Comparing it with the known equation of wave,

  y(x,t)= y_osin(kx+\omega t)

Assuming Amplitude to be in "cm" and "wavevector and angular frequency" be in SI units

a). We get amplitude of the wave as

  y_o= 3.96cm(ANS)

b). Wave-vector of the wave equation as k= 0.0444 \pi

Then wavelength of the wave will be

  \lambda =\frac{2 \pi}{k}= \frac{2 \pi }{0.0444 \pi}= 45.045 m(ANS)

c). Similarly angular frequency will be

\omega= 7.89 \pi

Then frequency will be

f= \frac{\omega}{2 \pi}= \frac{7.89 \pi}{2 \pi }= 3.945 Hz(ANS)

d). Speed of the wave is given by

v= f\lambda=3.945 \times 45.045= 177.70 m/s(ANS)

e). Since the coefficient of the wavevector is positive hence this wave propagates in "negative -x" direction.

f). Differentiating given tranverse displacement equation,

V= \frac{dy}{dt}= (3.96cm)(7.89 \pi)cos(0.0444 \pi x+7.89 \pi t)

Then maximum transverse speed will be

V_{max}= 3.96(7.89 \pi) cm/s= 98.107 cm/s(ANS)

f). Using given values in given transverse wave equation,

  y(x,t)= 3.96sin(0.0444\pi x+7.89\pi t)

y(x,t)= 3.96sin(0.0444\pi (0.0105)+7.89\pi (0.843))

y(0.0105,0.843)= (3.96cm)cos(0.0004662 \pi+6.6512\pi )= (3.96cm)cos(6.6516\pi )

y(0.0105,0.843)= -1.815cm(ANS)